# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ 4.2 Microphone subset selection
As explained in Section 4.1, microphone subset selection is an essential component in SE systems that need to handle distributed microphones. To select effective microphones, the conventional method implemented in the baseline system relies only on a single audio feature, i.e., EV (Section 4.2.1), which may not be robust enough when dealing with diverse recording scenarios, such as in this CHiME-8 challenge. We, therefore, propose leveraging multiple audio features to enhance robustness against environmental changes. We also propose imposing a lower bound on the number of microphones to be selected based on the assumption that the SE system can benefit from the spatial diversity of distributed microphones unless the number of microphones exceeds a certain threshold. We describe in Section 4.2.2 how these ideas are integrated into a single module.
@@ 4.2.1 Conventional method based on envelope variance
To select the top $K$ most effective microphones out of $M$ microphones, where $K$ ($\leq M$) is specified by the user, the conventional microphone subset selection module first calculates a signal quality measure called an EV [59] of the input signal for each microphone. The EV is a signal-level measure that is expected to have lower values when the input signal is more degraded by reverberation and background noise, and higher values when the signal is less degraded. For a definition and detailed description of the EV, we refer readers to [59]. The module then selects the $K$ microphones with the highest EV scores. In the baseline system, $K$ is set to $K=\lceil 0.8\times M\rceil$ across all datasets and remains fixed, not varying depending on the dataset.
@@ 4.2.2 Proposed method based on EV and speech clarity index $C_{50}$
First, we discuss the second point regarding the minimum number of microphones to select. In the SE system, we assume that using at least $K_{\min}$ microphones always improves the quality of the enhanced signal in the SE system. This is because using more microphones generally leads to a greater gain in SNR of the beamforming output signal, motivating us to select as many microphones as possible. On the other hand, the improvement in SNR by beamforming tends to saturate as the number of microphones increases. Moreover, selecting ineffective microphones, such as those with very low SNR, could negatively impact the beamforming output signal. The same argument for beamforming also applies to the TF mask estimation using GSS. Based on these observations, we hypothesize that discarding only ineffective microphones is more advantageous than selecting a small number of highly effective microphones. We therefore impose a lower bound, $K_{\min}$, on the minimum number of microphones to select and pass to the subsequent module. The value of $K_{\min}=15$ is empirically determined using the development set. For experimental results, see A.2.2.
@@
Let $\mathcal{K}_{\mathrm{EV}}\subseteq\{1,\ldots,M\}$ and $\mathcal{K}_{C_{50}}\subseteq\{1,\ldots,M\}$ be the sets of the top $K_{1}=\lceil 0.65\times M\rceil$ microphones ranked by EV and $C_{50}$, respectively, where we determine the value $K_{1}$ empirically. Let $\mathcal{K}=\mathcal{K}_{\mathrm{EV}}\cap\mathcal{K}_{C_{50}}$ denote the intersection of the two sets. According to the definition, $|\mathcal{K}|\leq K_{1}=|\mathcal{K}_{\mathrm{EV}}|=|\mathcal{K}_{C_{50}}|$. We propose the following algorithm for selecting a set of microphones to pass to the next module in the SE system:666 The algorithm presented in our workshop paper [21] contained a minor error in the fallback mechanism, which we correct in this paper. We did not devise this fallback mechanism, and we just used it in the CHiME-8 challenge.
@@
    If $M>K_{\min}$ and $|\mathcal{K}|\geq K_{\min}$, then we select $\mathcal{K}=\mathcal{K}_{\mathrm{EV}}\cap\mathcal{K}_{C_{50}}$.
@@
4.3 MIMO beamforming, reference microphone selection, and postfiltering
This section describes in detail three modules in Figure 7: MIMO beamforming (Sections 4.3.2 and 4.3.3), reference microphone selection (Section 4.3.4), and postfiltering (Section 4.3.5). While the baseline system implements R1-MWF [60] for the MIMO beamforming, we propose replacing it with Spatial-Prediction MWF (SP-MWF) [31, 30] as described in Section 4.3.3. As we will see, this change to the MIMO beamformer only affects the frequency-wise amplitude of the MIMO beamforming output signal, but it leads to the selection of a different reference microphone in the reference microphone selection module, which may be crucial in distributed microphone scenarios. We discuss this in Section 4.3.6.
@@ 4.3.1 Signal model
With a slight abuse of notation, let $\mathbf{y}(f,t)\in\mathbb{C}^{K}$ with frequency bin index $f$ and time frame index $t$ be the STFT representation of the $K$-channel dereverberated signal corresponding to utterance segment $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$ in the time domain. Suppose that the input signal $\mathbf{y}(f,t)$ is the superposition of the source image of the target source $s$, denoted as $\mathbf{x}_{\mathrm{s}}(f,t)$, and a noise-plus-interference component (hereafter simply referred to as noise for short), denoted as $\mathbf{x}_{\mathrm{n}}(f,t)$:
@@
is as close as possible to the target source component $\mathbf{x}_{\mathrm{s}}(f,t)$, where H represents the Hermitian transpose. To estimate $\mathbf{W}(f)$, the beamforming module exploits two TF masks provided by the GSS module: one for the target-plus-noise bins, denoted as $\alpha_{\mathrm{s}}(f,t)$,777 Here, strictly speaking, we cannot isolate the ambient noise from the target speech, and $s$ should be understood as target-plus-noise. and the other for the noise-only bins, denoted as $\alpha_{\mathrm{n}}(f,t)$. When considering the TF masks, we assume that the noise signal is present in all TF bins. In what follows, we introduce several MIMO beamformers and how they are obtained using the TF masks. We often omit the indices $f$ and $t$ to simplify the notations without causing confusion.
To explain MIMO beamformers, we introduce some notation. Let $\mathbf{z}(f)\in\mathbb{C}^{K}$ be the acoustic transfer function (ATF) or steering vector of the target source $s$ to the $K$ microphones, and let
@@
where $\mathcal{T}$ denotes the set of time frames corresponding to the considered segment, $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$, in the time domain. In the ideal case, $\mathbf{R}_{\mathrm{s}}$ is of rank 1 and has the following relation:
@@
where $P_{\mathrm{s}}\in\mathbb{R}_{\geq 0}$ is the power spectral density of the target source. Note, however, that $\mathbf{R}_{\mathrm{s}}$ estimated as in Eq. (20) is generally not of rank 1.
@@ 4.3.2 MVDR and R1-MWF
The standard MVDR beamformer with reference microphone $r\in\{1,\ldots,K\}$ is given as
$\mathbf{w}_{r}=\frac{\mathbf{R}_{\mathrm{n}}^{-1}\mathbf{z}}{\mathbf{z}^{\mathsf{H}}\mathbf{R}_{\mathrm{n}}^{-1}\mathbf{z}}z_{r}^{*},$ (22)
where $z_{r}\in\mathbb{C}$ denotes the $r$th element of $\mathbf{z}$ and $z_{r}^{*}$ is its complex conjugate. This form of the MVDR beamformer estimates the target source as observed at the $r$th microphone. If the relation in Eq. (21) holds, Eq. (22) is also expressed as [30, 60]:
@@
$\operatorname*{minimize}_{\mathbf{w}_{r}}$ $\mathbb{E}[\,|\mathbf{w}_{r}^{\top}\mathbf{x}_{\mathrm{s}}-\mathbf{u}_{r}^{\mathsf{H}}\mathbf{x}_{\mathrm{s}}|^{2}\,]+\gamma\mathbb{E}[\,|\mathbf{w}_{r}^{\mathsf{H}}\mathbf{x}_{\mathrm{n}}|^{2}\,].$ (24)
@@ 4.3.3 Spatial-prediction MWF (SP-MWF)
The equivalence under $\operatorname{rank}\mathbf{R}_{\mathrm{s}}=1$ or $\mathbf{R}_{\mathrm{s}}=P_{\mathrm{s}}\mathbf{z}\mathbf{z}^{\mathsf{H}}$ can be confirmed by checking that the second term of the denominator in Eq. (27) is equivalent to that of R1-MWF in Eq. (26), as follows:
@@
Then, substituting $\mathbf{z}$ at Eq. (30) into the MVDR formula of Eq. (22), we obtain
@@
As we mentioned, R1-MWF and SP-MWF are identical when $\operatorname{rank}\mathbf{R}_{\mathrm{s}}=1$ and proportional even when $\operatorname{rank}\mathbf{R}_{\mathrm{s}}\geq 2$. Although they differ only in the norm $\|\mathbf{w}_{r}(f)\|_{2}$ at each frequency bin, this difference may significantly affect the output of the reference microphone selection module. We will return to this topic in Section 4.3.6 after reviewing the reference microphone selection and postfiltering modules in Sections 4.3.4 and 4.3.5, respectively.
@@ 4.3.4 Reference microphone selection
which is then passed to the final postfiltering module.
4.3.5 Postfiltering
The postfiltering module in the baseline system receives $\hat{x}_{\mathrm{s},r_{\star}}(f,t)$ in Eq. (33) and applies the BAN postfilter [63] to it:
@@ 4.3.5 Postfiltering
Note that, unlike the baseline system, our SE system does not apply the BAN postfilter, as it is not effective when used in conjunction with SP-MWF. We refer readers to A.2.1 for the corresponding ablation study.
After that, the postfiltering module applies TF masking to further enhance the signal:
@@
Here, the TF mask $\alpha_{\mathrm{s}}(f,t)$ is floored to prevent excessive enhancement that could cause speech distortion. The flooring coefficient $0\leq\delta\leq 1$ is set to $-9$ (dB), or $\delta\approx 0.355$, in the default setting.
@@
In the reference microphone selection rule at Eq. (32), the norm of the filter, $\|\mathbf{w}_{r}(f)\|_{2}^{2}$, determines the frequency-wise weight of the separated components when calculating the output SNR. As a result, the selected microphone $r_{\star}$ may frequently differ depending on whether SP-MWF or R1-MWF is used for the MIMO beamformer $\mathbf{w}_{r}(f)$. In a distributed microphone scenario, such as the one considered in this CHiME challenge, selecting an effective microphone, e.g., microphone near the target source with a high SNR, can substantially improve the quality of the separated signal, compared to selecting an ineffective microphone. Although determining which beamformer is theoretically more effective for reference microphone selection remains an open question, we empirically demonstrate in A.2.1 that SP-MWF outperforms R1-MWF on the development set of this challenge.
It is worth noting that SP-MWF and R1-MWF differ only in the reference microphone selection if we apply the BAN postfilter to both. To see this, let us consider the case where the same reference microphone $r^{\star}$ is selected for both beamformers, and the BAN postfilter is applied in the postfiltering module. In this case, the output of the SE system is expressed as
@@
which is independent of the scale of the beamformer $\mathbf{w}_{r_{\star}}(f)$ due to the BAN postfilter. This implies that when the BAN postfilter is applied, the performance difference between SP-MWF and R1-MWF arises solely from the difference in reference microphone selection. We will demonstrate this in an ablation study in A.2.1.
@@ 4.4 Chunk-wise SE diarization
The SE processing we used within the diarization module to obtain better speaker embeddings, as explained in Section 3.1, is shown in Fig. 8. It differs slightly from the SE processing explained above and shown in Fig. 7. There are two main differences, GSS is performed chunk-wise, i.e., given the local speaker activity computed by the EEND model from each chunk, instead of utterance-wise, relying on estimated utterance boundaries as described in Eq. (4). Note that no context was used (i.e. $\Delta=0$), making it different from the utterance-wise SE approach. The other difference is that we perform GSS $M$-times, each time given the chunk activity estimated from a different microphone. This results in having $M$ TF-masks. Then, MIMO beamforming and reference channel selection (without post-filtering) are performed $M$ times independently, each time with a different mask. We used the same hyperparameters as those used by the proposed SE system $\operatorname{SE}(\cdot)$, except for $\delta=1$ in Eq. (35), meaning that no masking was applied to the beamforming output signal $\hat{x}_{\mathrm{s},r_{\star}}$ given by Eq. (33).
The output of the chunkwise SE module consists thus of $M$ enhanced signals, which allows the computation of enhanced speaker embeddings for each local speaker and each microphone. We analyze the effect of SE on diarization and speaker counting in Appendices A.1.1 and A.1.2.

### v2

@@ 4.2 Microphone subset selection
As explained in Section 4.1, microphone subset selection is an essential component in SE systems that need to handle distributed microphones. To select effective microphones, the conventional method implemented in the baseline system relies only on a single audio feature, i.e., EV (Section 4.2.1), which may not be robust enough to deal with diverse recording scenarios, such as those in this CHiME-8 challenge. We therefore propose leveraging multiple audio features to enhance robustness against environmental changes. We also propose imposing a lower bound on the number of microphones to be selected based on the assumption that the SE system can benefit from the spatial diversity of distributed microphones unless the number of microphones exceeds a certain threshold. Section 4.2.2 explains how these ideas are integrated into a single module.
@@ 4.2.1 Conventional method based on envelope variance
To select the top $K$ most effective microphones out of $M$ microphones, where $K$ ($\leq M$) is specified by the user, the conventional microphone subset selection module first calculates a signal quality measure called an EV [62] of the input signal for each microphone. The EV is a signal-level measure that is expected to have lower values when the input signal is more degraded by reverberation and background noise, and higher values when the signal is less degraded. For the definition and detailed description of the EV, we refer readers to [62]. The module then selects the $K$ microphones with the highest EV scores. In the baseline system, $K$ is set to $K=\lceil 0.8\cdot M\rceil$ across all datasets and remains fixed, i.e., not varying depending on the dataset.
@@ 4.2.2 Proposed method based on EV and speech clarity index $C_{50}$
First, we discuss the second point regarding the minimum number of microphones to select. In the SE system, we assume that using at least $K_{\min}$ microphones always improves the quality of the enhanced signal in the SE system. This is because using more microphones generally leads to a greater gain in the SNR of the beamforming output signal, motivating us to select as many microphones as possible. On the other hand, the SNR improvement by beamforming tends to saturate as the number of microphones increases. Moreover, selecting ineffective microphones, such as those with a very low SNR, could negatively impact the beamforming output signal. The same argument for beamforming also applies to the TF mask estimation using GSS. In the basis of on these observations, we hypothesize that discarding only ineffective microphones is more advantageous than selecting a small number of highly effective microphones. We therefore impose a lower bound, $K_{\min}$, on the minimum number of microphones to select and pass to the subsequent module. The value of $K_{\min}=15$ is empirically determined using the development set. For experimental results, see A.2.2.
@@
Let $\mathcal{K}_{\mathrm{EV}}\subseteq\{1,\ldots,M\}$ and $\mathcal{K}_{C_{50}}\subseteq\{1,\ldots,M\}$ be the sets of the top $K_{1}=\lceil 0.65\cdot M\rceil$ microphones ranked by EV and $C_{50}$, respectively, where we determine the value $K_{1}$ empirically. Let $\mathcal{K}=\mathcal{K}_{\mathrm{EV}}\cap\mathcal{K}_{C_{50}}$ denote the intersection of the two sets. According to the definition, $|\mathcal{K}|\leq K_{1}=|\mathcal{K}_{\mathrm{EV}}|=|\mathcal{K}_{C_{50}}|$. We propose the following algorithm for selecting a set of microphones to pass to the next module in the SE system:666 The algorithm presented in our workshop paper [21] contained a minor error in the fallback mechanism, which we have corrected in this paper. Note that we did not revise this fallback mechanism and we just used it in the CHiME-8 challenge.
@@
    If $M>K_{\min}$ and $|\mathcal{K}|\geq K_{\min}$, we select $\mathcal{K}=\mathcal{K}_{\mathrm{EV}}\cap\mathcal{K}_{C_{50}}$.
@@
4.3 MIMO beamforming, reference microphone selection, and post-filtering
This section describes in detail the three modules shown in Figure 7: MIMO beamforming (Sections 4.3.2 and 4.3.3), reference microphone selection (Section 4.3.4), and post-filtering (Section 4.3.5). While the baseline system implements R1-MWF [63] for the MIMO beamforming, we propose replacing it with Spatial-Prediction MWF (SP-MWF) [31, 30] as described in Section 4.3.3. As we will see, although this change to the MIMO beamformer only affects the frequency-wise amplitude of the MIMO beamforming output signal, it leads to the selection of a different reference microphone in the reference microphone selection module, which may be crucial in distributed microphone scenarios. We discuss this in more detail in Section 4.3.6.
@@ 4.3.1 Signal model
With a slight abuse of notation, let $\mathbf{y}(f,t)\in\mathbb{C}^{K}$ with frequency bin index $f$ and time frame index $t$ be the STFT representation of the $K$-channel dereverberated signal corresponding to utterance segment $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$ in the time domain. Suppose that the input signal $\mathbf{y}(f,t)$ is the superposition of the source image of the target source $s$, denoted as $\mathbf{x}_{\mathrm{s}}(f,t)$, and a noise-plus-interference component (hereafter noise), denoted as $\mathbf{x}_{\mathrm{n}}(f,t)$
@@
is as close as possible to the target source component $\mathbf{x}_{\mathrm{s}}(f,t)$, where H represents the Hermitian transpose. To estimate $\mathbf{W}(f)$, the beamforming module exploits two TF masks provided by the GSS module: one for the target-plus-noise bins, denoted as $\alpha_{\mathrm{s}}(f,t)$,777 Here, strictly speaking, we cannot isolate the ambient noise from the target speech, and $s$ should be understood as target-plus-noise. and the other for the noise-only bins, denoted as $\alpha_{\mathrm{n}}(f,t)$. When considering the TF masks, we assume that the noise signal is present in all TF bins. In what follows, we introduce several MIMO beamformers and explain how they are obtained using the TF masks. We often omit the indices $f$ and $t$ to simplify the notations without causing confusion.
To explain MIMO beamformers, we first introduce some notation. Let $\mathbf{z}(f)\in\mathbb{C}^{K}$ be the acoustic transfer function (ATF) or steering vector of the target source $s$ to the $K$ microphones, and let
@@
where $\mathcal{T}$ denotes the set of time frames corresponding to the considered segment, $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$, in the time domain. In the ideal case, $\mathbf{R}_{\mathrm{s}}$ is of rank 1 and has the following relation [69]:
@@
where $P_{\mathrm{s}}\in\mathbb{R}_{\geq 0}$ is the power spectral density of the target source. Note, however, that in reverberant conditions where the window size of the STFT is shorter than the reverberation time, $\mathbf{R}_{\mathrm{s}}$ can have a rank greater than 1. Note also that $\mathbf{R}_{\mathrm{s}}$ estimated as in Eq. (20) is generally not of rank 1.
@@ 4.3.2 MVDR and R1-MWF
The MVDR beamformer with reference microphone $r\in\{1,\ldots,K\}$ is given as [69]
$\mathbf{w}_{r}$ $=\operatorname*{argmin}_{\mathbf{w}_{r}}\{\mathbf{w}_{r}^{\mathsf{H}}\mathbf{R}_{\mathrm{n}}\mathbf{w}_{r}\mid\mathbf{w}_{r}^{\mathsf{H}}\mathbf{z}=z_{r}\}$
$=\frac{\mathbf{R}_{\mathrm{n}}^{-1}\mathbf{z}}{\mathbf{z}^{\mathsf{H}}\mathbf{R}_{\mathrm{n}}^{-1}\mathbf{z}}z_{r}^{*},$ (22)
where $z_{r}\in\mathbb{C}$ denotes the $r$th element of $\mathbf{z}$ and $z_{r}^{*}$ is its complex conjugate. This form of the MVDR beamformer estimates the target source as observed at the $r$th microphone. If the relation in Eq. (21) holds, Eq. (22) is also expressed as [30, 63]
@@
$\operatorname*{minimize}_{\mathbf{w}_{r}}$ $\mathbb{E}[\,|\mathbf{w}_{r}^{\mathsf{H}}\mathbf{x}_{\mathrm{s}}-\mathbf{u}_{r}^{\top}\mathbf{x}_{\mathrm{s}}|^{2}\,]+\gamma\mathbb{E}[\,|\mathbf{w}_{r}^{\mathsf{H}}\mathbf{x}_{\mathrm{n}}|^{2}\,].$ (24)
@@ 4.3.3 Spatial-prediction MWF (SP-MWF)
The equivalence under $\operatorname{rank}\mathbf{R}_{\mathrm{s}}=1$ or $\mathbf{R}_{\mathrm{s}}=P_{\mathrm{s}}\mathbf{z}\mathbf{z}^{\mathsf{H}}$ can be confirmed by checking that the second term of the denominator in Eq. (27) is equivalent to that of R1-MWF in Eq. (26), as
@@
Then, substituting $\mathbf{z}$ in Eq. (30) into the MVDR formula of Eq. (22), we obtain
@@
As we mentioned, R1-MWF and SP-MWF are identical when $\operatorname{rank}\mathbf{R}_{\mathrm{s}}=1$ and proportional even when $\operatorname{rank}\mathbf{R}_{\mathrm{s}}\geq 2$. Although they differ only in the norm $\|\mathbf{w}_{r}(f)\|_{2}$ at each frequency bin, this difference may significantly affect the output of the reference microphone selection module. We will return to this topic in Section 4.3.6 after reviewing the reference microphone selection and post-filtering modules in Sections 4.3.4 and 4.3.5, respectively.
@@ 4.3.4 Reference microphone selection
which is then passed to the final post-filtering module.
4.3.5 post-filtering
The post-filtering module in the baseline system receives $\hat{x}_{\mathrm{s},r_{\star}}(f,t)$ in Eq. (33) and applies the BAN post-filter [66] to it:
@@
Note that, unlike the baseline system, our SE system does not apply the BAN post-filter, as it is not effective when used in conjunction with SP-MWF. We refer readers to A.2.1 for the corresponding ablation study.
After that, the post-filtering module applies TF masking to further enhance the signal:
@@
Here, the TF mask $\alpha_{\mathrm{s}}(f,t)$ is floored to prevent excessive enhancement that could cause speech distortion. The flooring coefficient $0\leq\delta\leq 1$ is set to $-9$ (dB), or $\delta\approx 0.355$, as a default.
@@
In the reference microphone selection rule in Eq. (32), the norm of the filter, $\|\mathbf{w}_{r}(f)\|_{2}^{2}$, determines the frequency-wise weight of the separated components when calculating the output SNR. As a result, the selected microphone $r_{\star}$ may frequently differ depending on whether SP-MWF or R1-MWF is used for the MIMO beamformer $\mathbf{w}_{r}(f)$. In a distributed microphone scenario, such as the one considered in this CHiME challenge, selecting an effective microphone, e.g., a microphone near the target source with a high SNR, can substantially improve the quality of the separated signal compared to selecting an ineffective microphone. Although determining which beamformer is theoretically more effective for reference microphone selection remains an open question, we empirically demonstrate in A.2.1 that SP-MWF outperforms R1-MWF on the development set of this challenge.
It is worth noting that SP-MWF and R1-MWF differ only in the reference microphone selection if we apply the BAN post-filter to both. To see this, let us consider the case where the same reference microphone $r^{\star}$ is selected for both beamformers, and the BAN post-filter is applied in the post-filtering module. In this case, the output of the SE system is expressed as
@@
which is independent of the scale of the beamformer $\mathbf{w}_{r_{\star}}(f)$ due to the BAN post-filter. This implies that when the BAN post-filter is applied, the performance difference between SP-MWF and R1-MWF arises solely from the difference in reference microphone selection. We will demonstrate this in the ablation study in A.2.1.
@@ 4.4 Chunk-wise SE diarization
The SE processing we used within the diarization module to obtain better speaker embeddings, as explained in Section 3.1, is shown in Figure 8. It differs slightly from the SE processing explained above and shown in Figure 7. There are two main differences, GSS is performed chunk-wise, i.e., it is given the local speaker activity computed by the EEND model from each chunk, instead of utterance-wise, and therefore relies on estimated utterance boundaries as described in Eq. (4). Note that no context is used (i.e. $\Delta=0$), making it different from the utterance-wise SE approach. The other difference is that we perform GSS $M$-times, which each time given the chunk activity estimated from a different microphone. This results in having $M$ TF-masks. Then, MIMO beamforming and reference channel selection (without post-filtering) are performed $M$ times independently, each time with a different mask. We utilized the same hyperparameters as those used by the proposed SE system $\operatorname{SE}(\cdot)$, except for $\delta=1$ in Eq. (35), meaning that no masking was applied to the beamforming output signal $\hat{x}_{\mathrm{s},r_{\star}}$ given by Eq. (33).
The output of the chunkwise SE module thus consists of $M$ enhanced signals, which allows the computation of enhanced speaker embeddings for each local speaker and each microphone. We analyze the effect of SE on diarization and speaker counting in Appendices A.1.1 and A.1.2.